# Understanding Stellar Mass-Metallicity and Size Relations in Simulated Ultra-Faint Dwarf Galaxies

### v1

@@ 3.3.4 Gradient in stellar metallicity
The discovery of member stars in the outskirts of observed UFDs raises an intriguing question: Is the metallicity gradient, where metal-poor stars are found farther from the center than metal-rich stars, also present in UFDs, similar to what is observed in more massive dwarf galaxies (e.g., Harbeck et al., 2001; Kirby et al., 2011; Fu et al., 2024)? Chiti et al. (2021) proposed the presence of a metallicity gradient in Tucana II, noting the detection of distant stars with $\rm[Fe/H]\approx-3.02$ at distances greater than $\gtrsim 9r_{\rm h}$. Furthermore, Tarumi et al. (2021) demonstrated through simulations that metallicity gradient in a galaxy with a virial mass of $M_{\rm vir}=2.5\times 10^{8}{\,M_{\odot}}$ at 3 Gyr can result from mergers between dwarf galaxies with different metallicities, where the metal-rich central galaxy stays intact and the metal-poor infalling galaxy is disrupted, spreading its stars to the outer regions.
@@
On the other hand, not all simulated UFDs in our study show a significant metallicity gradient. Fig. 18 illustrates the stellar metallicities as a function of 3D distances from the center of each UFD analog at $z=0$. Stars are color-coded according to their progenitor halos to identify their origins. For instance, in the case of Halo1, there are seven extended stars lying in $r\gtrsim 3\ r_{\rm h}$, which have higher metallicities than those lying in $r<3\ r_{\rm h}$, resulting in no clear metallicity gradient between these two populations. Furthermore, as discussed in Section 3.2.2, low-metallicity stars with $\rm[Fe/H]<-4$, shown in Halo2--Halo5, mainly result from external metal enrichment. These stars do not predominantly reside in the outer regions and constitute only a small portion of the total population, thus not producing the overall metallicity gradient.
@@
Notably, Halo5 exhibits high-metallicity stars at larger radii ($\gtrsim 1$ kpc), predominantly originating from a single progenitor halo, prog5 (purple) in Fig. 18. Star formation in the progenitor prog5 of Halo5 began relatively late, around $z\lesssim 8$ (see also Fig. 11), resulting in stars with metallicities ranging between $-3\lesssim\rm[Fe/H]\lesssim-1.5$. These stars, spanning a wide metallicity range, are distributed over an extensive spatial range, up to $\sim$10 kpc at $z=0$. This dispersion is a consequence of a late merger at $z\approx 0.7$, leading to the formation of extended structures. We point out that if current observational capabilities are limited to the central regions ($\lesssim 3\ r_{\rm h}$), as indicated by the vertical line in Fig. 18, our simulations suggest that the metallicity properties will be primarily determined by stars formed from $1-3$ progenitor halos located in the inner region at $z=0$. For instance, in the case of Halo2 within this region, stars from three different progenitors (prog1, prog2, and prog4) contribute to the metallicity analysis.
@@
The generation of the metallicity distribution as a function of distance in our simulated UFDs involves several steps: (1) Stars form from the central star-forming gas cloud, with the MDF depicted in the top panel of Fig. 19. (2) SN feedback from these stars clears the central region of the progenitor halo, causing the shock front to move outward beyond $0.1R_{\rm vir}$. (3) Although gas within $0.1R_{\rm vir}$ is entirely ejected, stars continue to form from the gas compressed by the shock front, which has been enriched by previous SN explosions, resulting in relatively high-metallicity stars ($\rm[Fe/H]\geq-2$). This process leads to the formation of relatively high-metallicity stars far from the center. In addition, due to the shallow gravitational potential well of the simulated UFDs, the central star-forming region is highly susceptible to stellar feedback, which makes sustained star formation in this region difficult. Consequently, this leads to a lack of a metallicity gradient.
Observational efforts to investigate the existence of a metallicity gradient over large distances are an ongoing area of research using high-resolution spectroscopy (e.g., Chiti et al., 2021; Sestito et al., 2023; Waller et al., 2023). Recently, Sestito et al. (2023) discovered five new member stars on the outskirts of Ursa Minor (one star at $\sim 12\ r_{\rm h}$ and four stars at $\sim 5.2-6.3\ r_{\rm h}$) using an updated identification technique proposed by Jensen et al. (2024). They reported that the metallicity of the most distant stars at $\sim 12\ r_{\rm h}$ is $\rm[Fe/H]\approx-2.09$, which is higher than the average value of $\rm[Fe/H]\approx-2.65$ for the four stars at $\sim 5.2-6.3\ r_{\rm h}$, although still lower than the inner region ($\lesssim 3\ r_{\rm h}$). Furthermore, Waller et al. (2023) reported the metallicity of newly discovered stars in three UFD galaxies: Coma Berenices, Bootes I, and Ursa Major I. Among them, they introduced two new stars located at $2.5\ r_{\rm h}$ of Coma Berenices, which exhibit the same metallicity range as the previously studied inner stars, with $-3<\rm[Fe/H]<-2$.
@@ 4 DISCUSSION AND CONCLUSION
In this study, we examine the physical properties, including stellar metallicity and galaxy size (i.e., half-light radius), of Ultra-Faint Dwarfs (UFDs), the faintest galaxy systems in the Universe ($L_{V}\gtrsim-7.7$) (e.g., Simon, 2019). Despite achieving high mass resolution ($m_{\rm gas}\lesssim 100{\,M_{\odot}}$), recent theoretical work predicts lower average stellar metallicity ($\rm-4\lesssim\langle[Fe/H]\rangle\lesssim-3$) compared to observed values in UFDs ($\rm-3\lesssim\langle[Fe/H]\rangle\lesssim-2$) and larger half-light radii ($r_{\rm h}>100$ pc) than those observed. To address this discrepancy in stellar metallicity and galaxy size between observations and theoretical predictions, we conduct a series of high-resolution hydrodynamic simulations ($m_{\rm gas}\approx 60{\,M_{\odot}},m_{\rm DM}\approx 370{\,M_{\odot}}$) on UFD analogs with virial masses $M_{\rm vir}\approx 2\times 10^{8}{\,M_{\odot}}-1.3\times 10^{9}{\,M_{\odot}}$ and stellar masses ranging from $M_{\star}\sim 10^{3}{\,M_{\odot}}$ to a few $10^{4}{\,M_{\odot}}$. In addition, we adopt an individual sampling method to extract massive stars for supernovae (SNe) explosions from a given initial mass function (IMF), allowing us to address the discrete nature of SNe.
@@
      The derived half-light radii of our simulated UFD analogs range from $r_{\rm h}\approx 89\ \rm pc$ ($M_{\star}=2.4\times 10^{3}\ ~{}\mbox{M}_{\odot}$) to $r_{\rm h}\approx 1115\ \rm pc$ ($M_{\star}=2.1\times 10^{4}\ ~{}\mbox{M}_{\odot}$) based on the direct calculation of star particles. To ensure consistency between simulations and observations, we employ the most commonly used observational method, which involves fitting a 2-D single exponential density profile to derive spatial structural parameters. This fitting method results in either similar values or a reduction in $r_{\rm h}$ by 20%, 30%, and 47% for Halo2, Halo3, and Halo5, respectively, compared to the direct method.
@@
      When utilizing the fitting method, accounting for background stars is crucial to accurately determine the half-light radius. Without including background stars, relying solely on stellar particles from simulations makes it challenging to obtain a best fit that captures both the inner and outer profiles, often resulting in an abnormally large half-light radius.
@@
      We find extended structures of varying degrees in all halos of our UFD analogs. Single exponential component fitting methods effectively describe the inner profile but struggle with extended outer structures. To address this issue, we adopt the recently proposed two-component stellar density profile (Jensen et al., 2024), which successfully captures both the inner and outer components in our simulated UFDs, allowing for more accurate estimation of structural parameters.
@@
As mentioned earlier, the presence and frequency of metal-rich stars provide vital insights into the environments of star-forming regions, a point also highlighted in other studies. In a recent study by Sandford et al. (2024), the authors offer an analytical prediction of the galactic evolution of the UFD galaxy Eridanus II (Eri II), suggesting that the abundance of metal-rich stars is a key indicator of various physical processes. They specifically assessed the influence of Type Ia SNe, which promote the formation of metal-rich stars by increasing iron yields, examining their effects on both the MDF and the $\rm[Mg/Fe]-[Fe/H]$ space. Similarly, Ting & Ji (2024) showed that analyzing the $\rm[\alpha/Fe]-[Fe/H]$ space can reveal whether dwarf galaxies experienced bursty star formation. However, in our simulations, Type Ia SNe are so rare that they had little effect, and the “knee” feature in the $\rm[\alpha/Fe]-[Fe/H]$ plane was not observed. This low frequency of Type Ia SNe is mainly due to the small stellar mass of our UFD analogs. Nevertheless, given that Type Ia SNe are not well constrained (e.g., Dubay et al., 2024; Alexander & Vincenzo, 2024), it would be interesting to examine whether the frequent Type Ia SNe depending on the delay-time to be triggered will change the MDF and the amount of high-metallicity stars in UFDs.
@@
Appendix A Resultant MDF of the UFD analogs with different SN energy
[…]

### v2

@@ 3.3.4 Gradient in stellar metallicity
The discovery of member stars in the outskirts of observed UFDs raises an intriguing question: Is the metallicity gradient, where metal-poor stars are found farther from the center than metal-rich stars, also present in UFDs, similar to what is observed in more massive dwarf galaxies (e.g., Harbeck et al., 2001; Kirby et al., 2011; Fu et al., 2024)? Chiti et al. (2021) proposed the presence of a metallicity gradient in Tucana II, noting the detection of distant stars with $\rm[Fe/H]\approx-3.02$ at distances greater than $\gtrsim 9\ r_{\rm h}$. Furthermore, Tarumi et al. (2021) demonstrated through simulations that a metallicity gradient in a galaxy with $M_{\rm vir}=2.5\times 10^{8}~{}\mbox{M}_{\odot}$ may arise from mergers between dwarf galaxies with different metallicities. In their scenario, the metal-rich, central galaxy remains intact while the metal-poor, infalling galaxy is disrupted, spreading its stars into the outer regions and thereby producing the gradient. Also, induced star formation during and after the merger further enhances the gradient.
@@
On the other hand, not all simulated UFDs in our study show such a metallicity gradient. Fig. 18 illustrates the stellar metallicities as a function of 2D elliptical radius from the center of each UFD analog at $z=0$. Stars are color-coded according to their progenitor halos to identify their origins. In the case of Halo1, for example, there are seven extended stars lying in $r\gtrsim 3\ r_{\rm h}$, which have higher metallicities than those lying in $r<3\ r_{\rm h}$, resulting in no clear metallicity gradient. Furthermore, as discussed in Section 3.2.2, low-metallicity stars with $\rm[Fe/H]<-4$ shown in Halo2--Halo5, are primarily result of external metal enrichment. However, these stars are not predominantly located in the outer regions and make up only a small fraction of the total stellar population; thus, they do not significantly contribute to the overall metallicity gradient.
@@
Notably, Halo5 exhibits high-metallicity stars at larger radii ($\gtrsim 1$ kpc), most of which originate from a single progenitor halo, prog5 (purple), as shown in Fig. 18. Star formation in the progenitor prog5 of Halo5 began relatively late, around $z\lesssim 8$ (see also Fig. 11), resulting in stars with metallicities ranging between $-3\lesssim\rm[Fe/H]\lesssim-1.5$. These stars, spanning a wide metallicity range, are distributed over an extensive spatial range, up to $\sim$10 kpc at $z=0$. This dispersion is a consequence of a late merger at $z\approx 0.7$, leading to the formation of extended structures. We point out that if current observational capabilities are limited to the central regions ($\lesssim 3\ r_{\rm h}$), as indicated by the vertical line in Fig. 18, our simulations suggest that the metallicity properties will be primarily determined by stars formed from $1-3$ progenitor halos located in the inner region at $z=0$. For instance, in the case of Halo2 within this region, stars from three different progenitors (prog1, prog2, and prog4) contribute to the metallicity analysis.
@@
The generation of the metallicity distribution as a function of distance in our simulated UFDs involves several steps: (1) Stars form from the central star-forming gas cloud, with the MDF depicted in the top panel of Fig. 19. (2) SNe feedback from these stars clears the central region of the progenitor halo, causing the shock front to move outward beyond $0.1R_{\rm vir}$. (3) Although gas within $0.1R_{\rm vir}$ is entirely ejected, stars continue to form from the gas compressed by the shock front, which has been enriched by previous SNe explosions, resulting in relatively high-metallicity stars ($\rm[Fe/H]\geq-2$). This process leads to the formation of relatively high-metallicity stars far from the center. In addition, due to the shallow gravitational potential well of the simulated UFDs, the central star-forming region is highly susceptible to stellar feedback, which makes sustained star formation in this region difficult. Consequently, this leads to a lack of a metallicity gradient. Observational efforts to investigate the existence of a metallicity gradient over large distances are an ongoing area of research using high-resolution spectroscopy (e.g., Chiti et al., 2021; Sestito et al., 2023; Waller et al., 2023). Recently, Sestito et al. (2023) discovered five new member stars on the outskirts of Ursa Minor (one star at $\sim 12\ r_{\rm h}$ and four stars at $\sim 5.2-6.3\ r_{\rm h}$) using an updated identification technique proposed by Jensen et al. (2024). They reported that the metallicity of the most distant stars at $\sim 12\ r_{\rm h}$ is $\rm[Fe/H]\approx-2.09$, which is higher than the average value of $\rm[Fe/H]\approx-2.65$ for the four stars at $\sim 5.2-6.3\ r_{\rm h}$, although still lower than metallicity of stars in the inner region ($\lesssim 3\ r_{\rm h}$). Furthermore, Waller et al. (2023) reported the metallicity of newly discovered stars in three UFD galaxies: Coma Berenices, Bootes I, and Ursa Major I. Among them, they introduced two new stars located at $2.5\ r_{\rm h}$ of Coma Berenices, which exhibit the same metallicity range as the previously studied inner stars, with $-3<\rm[Fe/H]<-2$.
@@ 4 DISCUSSION AND CONCLUSION
In this study, we examine the physical properties, including stellar metallicity and galaxy size (i.e., half-light radius), of Ultra-Faint Dwarfs (UFDs), the faintest galaxy systems in the Universe ($L_{V}\gtrsim-7.7$) (e.g., Simon, 2019). Despite achieving high mass resolution ($m_{\rm gas}\lesssim 100{\,M_{\odot}}$), recent theoretical work predicts lower average stellar metallicity ($\rm-4\lesssim\langle[Fe/H]\rangle\lesssim-3$) compared to observed values in UFDs ($\rm-3\lesssim\langle[Fe/H]\rangle\lesssim-2$) and larger half-light radii ($r_{\rm h}>100$ pc) than those observed. To address this discrepancy in stellar metallicity and galaxy size between observations and theoretical predictions, we conduct a series of high-resolution hydrodynamic simulations ($m_{\rm gas}\approx 60{\,M_{\odot}},m_{\rm DM}\approx 300{\,M_{\odot}}$) on UFD analogs with virial masses $M_{\rm vir}\approx 2\times 10^{8}{\,M_{\odot}}-1.3\times 10^{9}{\,M_{\odot}}$ and stellar masses ranging from $M_{\star}\sim 10^{3}{\,M_{\odot}}$ to a few $10^{4}{\,M_{\odot}}$. In addition, we adopt an individual sampling method to extract massive stars for supernovae (SNe) explosions from a given initial mass function (IMF), allowing us to address the discrete nature of SNe.
@@
      For size measurements of UFD analogs, to ensure consistency between simulations and observations, we use the most commonly adopted observational approach, which fits a two-dimensional single exponential surface density profile to derive spatial structural parameters. We also compute the half-light radius using the direct method commonly used in simulations, where the radius is defined as the distance enclosing half of the total stellar mass based on star particles.
      When using the fitting method, it is important to account for artificial background stars because the presence of extended star particles in UFD analogs can distort the surface density profile. This distortion may lead to an overestimation of the half-light radius if background stars are not properly considered. In fact, for UFD analogs, the half-light radius derived without background stars ($r_{\rm fit,\ (w/o)}$) is, on average, three times larger than that obtained with background stars ($r_{\rm fit,\ (w/)}$). Furthermore, $r_{\rm fit,\ (w/)}$ shows similar values compared to the those from the direct method ($r_{\rm dir}$). This highlights the necessity of accounting for artificial background stars to recover the true stellar distribution.
@@
      We find extended structures of varying degrees in all halos of our UFD analogs. Single exponential component fitting methods with background stars effectively describe the inner profile but still struggle with extended outer structures. To address this issue, we adopt the recently proposed two-component stellar density profile (Jensen et al., 2024), which successfully captures both the inner and outer components in our simulated UFDs, allowing for more accurate estimation of structural parameters.
@@
      We find extended structures of varying degrees in all halos of our UFD analogs. Single exponential component fitting methods with background stars effectively describe the inner profile but still struggle with extended outer structures. To address this issue, we adopt the recently proposed two-component stellar density profile (Jensen et al., 2024), which successfully captures both the inner and outer components in our simulated UFDs, allowing for more accurate estimation of structural parameters.
@@
We emphasize that we do not account for delay time for stars dying as SNe, as a means to compensate for the absence of RT. Otherwise, considering delay time without RT would artificially increase star formation and lead to overestimating both stellar mass and metallicity. A more accurate approach should consider delay time with radiative feedback, which will be addressed in future work. The work by Smith et al. (2021) found that incorporating RT could suppress the clustering of SNe, resulting in a less effective outflow rate compared to scenarios without RT, while maintaining the total number of SNe. Given this effect, accounting for RT would lead to a more dispersed distribution of SNe, potentially reducing the overall stellar metallicity. Consequently, a more realistic cosmological simulation with advanced feedback mechanisms would likely yield a lower average metallicity than our current results, suggesting that the stellar metallicity observed in this study may represent an upper limit. However, intriguingly, in such a case, the simulated UFD analogs would exhibit lower metallicity compared to the observed UFDs, therefore further widening the gap between observational data and theoretical models.
As mentioned earlier, the presence and frequency of metal-rich stars provide vital insights into the environments of star-forming regions, a point also highlighted in other studies. In a recent study by Sandford et al. (2024), the authors offer an analytical prediction of the galactic evolution of the UFD galaxy Eridanus II (Eri II), suggesting that the abundance of metal-rich stars is a key indicator of various physical processes. They specifically assessed the influence of Type Ia SNe, which promote the formation of metal-rich stars by increasing iron yields, examining their effects on both the MDF and the $\rm[Mg/Fe]-[Fe/H]$ space. Similarly, Ting & Ji (2024) showed that analyzing the $\rm[\alpha/Fe]-[Fe/H]$ space can reveal whether dwarf galaxies experienced bursty star formation. However, in our simulations, Type Ia SNe are so rare that they had little effect, and the “knee” feature in the $\rm[\alpha/Fe]-[Fe/H]$ plane was not observed. This low frequency of Type Ia SNe is mainly due to the small stellar mass of our UFD analogs. Nevertheless, given that Type Ia SNe are not well constrained (e.g., Dubay et al., 2024; Alexander & Vincenzo, 2024), it would be interesting to examine whether frequent Type Ia SNe, depending on the delay time distribution, alter the MDF and the abundance of high-metallicity stars in UFDs.
@@ acknowledgements
Appendix A Resultant MDF of the UFD analogs with different SN energy
[…]
Appendix B Alpha element abundances
We investigate the $\rm[Mg/Fe]\ vs\ [Fe/H]$ distributions of member stars in our UFD analogs. For comparison with observations, we select galaxies with absolute magnitudes fainter than $M_{\rm V}\sim-6.3$, which is similar to that of our most luminous UFD analog Halo5. These observed galaxies include Horologium I (Nagasawa et al., 2018), Reticulum II (Ji et al., 2016; Ji & Frebel, 2018), Bootes II (Ji et al., 2016; François et al., 2016), Tucana II (Chiti et al., 2018), Canes Venatici II (Vargas et al., 2013; François et al., 2016), Coma Berenices (Frebel et al., 2010; Vargas et al., 2013), Segue II (Kirby et al., 2013), Triangulum II (Kirby et al., 2017; Ji et al., 2019), Tucana III (Hansen et al., 2017), Ursa Major II (Frebel et al., 2010; Vargas et al., 2013), Grus I (Ji et al., 2019), Segue I (Vargas et al., 2013; Frebel et al., 2014), Leo IV (Simon et al., 2010; François et al., 2016), and Bootes I (Ishigaki et al., 2014; Gilmore et al., 2013). Although our data generally agree well with observations, the low-metallicity stars with $\rm[Fe/H]\lesssim-4$ exhibit extremely high $\rm[Mg/Fe]$ ratios. This is primarily due to the influence of Pop III SN, which yields a significantly higher abundance of alpha-elements than Pop II CCSNe.
Figure 22: $\rm[Mg/Fe]\ vs.[Fe/H]$ distribution for simulated UFD analogs, with stars represented as circles colored according to their progenitor halos, compared with observed data shown in grey.
Appendix C Two-component fitting results of surface brightness profile
Fig. 23 shows results of the two-component surface brightness fitting for other UFD analogs. We perform the same fitting method as represented in Section 3.3.
Figure 23: The surface brightness profiles of UFD analogs are modeled using a two-component approach. The actual density profile is shown as black-filled circles at binned radii. The inner component is depicted in red, while the outer component is illustrated in blue. A vertical line indicates the transition radius between the inner and outer components.